# Interpretability by Design for Efficient Multi-Objective Reinforcement Learning

### v1

@@ B.1 Benchmarks
To evaluate the performance of our proposed LLE-MORL method and compare it against existing baselines, we utilise a suite of continuous control benchmarks from the MO-Gymnasium library [Felten et al., 2023]. These environments are designed to test the ability of an agent to learn policies that effectively balance multiple, often conflicting objectives. The specific environments and their multi-objective reward formulations are detailed below:
@@
This task involves controlling a three-link Swimmer robot to swim forward quickly while minimizing energy expenses. The observation space $\mathcal{S}\subset\mathbb{R}^{8}$ includes joint angles and velocities, and the action space $\mathcal{A}\subset\mathbb{R}^{2}$ represents joint torques in $[-1,1]$. Let $x_{\text{before}}$ and $x_{\text{after}}$ be the $x$-coordinates of the centre of mass of swimmer before and after an action, $\Delta t$ be the time step, and $a_{j}$ be the $j$-th component of the action vector.
@@
In this task, a one-legged Hopper robot aims for forward locomotion and achieving height, with control costs and survival rewards influencing both objectives. The observation space $\mathcal{S}\subset\mathbb{R}^{11}$ includes joint states and torso position, and the action space $\mathcal{A}\subset\mathbb{R}^{3}$ represents joint torques in $[-1,1]$. Let $v_{x}=(x_{\text{after}}-x_{\text{before}})/{\Delta t}$ be the forward velocity of the agent along the $x$-axis, where $x_{\text{after}}$ and $x_{\text{before}}$ are x-positions of the torso. Let $h_{\text{jump}}=10\times(z_{\text{after}}-z_{\text{init}})$ be a measure of jumping height, where $z_{\text{after}}$ is the current z-position of the torso and $z_{\text{init}}$ is its initial z-position. Let $c_{\text{ctrl}}$ be the positive control cost, computed as $w_{\text{env\_ctrl}}\sum_{j}(a_{j})^{2}$, where $w_{\text{env\_ctrl}}$ is the environment control cost weight (typically $0.001$). Let $r_{\text{healthy}}$ be the health reward (typically $+1$ if the agent has not fallen). The reward vector $\mathbf{R}=[R_{1},R_{2}]$ is defined as:
@@
This task involves controlling a four-legged Ant robot. For the 2-objective version, the agent aims to achieve locomotion, with objectives based on $x$-velocity and $y$-velocity, both adjusted by control costs, survival rewards, and contact penalties. The observation space $\mathcal{S}\subset\mathbb{R}^{27}$ includes joint states, torso position, and contact forces, and the action space $\mathcal{A}\subset\mathbb{R}^{8}$ represents joint torques in $[-1,1]$. Let $v_{x}=(x_{\text{after}}-x_{\text{before}})/{\Delta t}$ be the forward velocity of the agent along the $x$-axis, where $x_{\text{after}}$ and $x_{\text{before}}$ are x-positions of the torso. Let $v_{y}=(y_{\text{after}}-y_{\text{before}})/{\Delta t}$ be the forward velocity of the agent along the $y$-axis, where $y_{\text{after}}$ and $y_{\text{before}}$ are y-positions of the torso. Let $c_{\text{ctrl}}$ be the positive control cost, computed as $w_{\text{env\_ctrl}}\sum_{j}(a_{j})^{2}$, where $w_{\text{env\_ctrl}}$ is the environment control cost weight (typically 0.05). Let $r_{\text{healthy}}$ be the health reward (typically $+1$ if the Ant is healthy). Let $p_{\text{contact}}$ be the positive contact penalty, which is used for penalising the Ant if the external contact forces are too large, computed as $w_{\text{env\_contact}}\sum_{k}(\text{force}_{k})^{2}$, where $w_{\text{env\_contact}}$ is the environment contact cost weight (typically $5\times 10^{-4}$). The reward vector $\mathbf{R}=[R_{1},R_{2}]$ is defined as:
@@
B.2 Training Details
All learning phases within our LLE-MORL algorithm, including the initial training of base policies, the directional retraining, and the final preference-aligned fine-tuning, utilize the Proximal Policy Optimization (PPO) algorithm [Schulman et al., 2017]. We employed a standard PPO implementation from the Stable Baselines3 library [Raffin et al., 2019]. The PPO parameters used across all training stages and benchmarks are detailed in Table 3.
@@ B.2 Training Details
  • Number of Initial Base Policies ($K$): The total count of base policies $\theta_{w_{i}}$, trained in the initialization stage. The corresponding $K$ initial scalarization weights $\{w_{i}\}_{i=1}^{K}$ are generated by evenly distributing them across the preference space (e.g., for 2D objectives, from $[1,0]$ to $[0,1]$ in $K$ steps).
  • Initialization Training Timesteps ($T_{\mathrm{init}}$): The number of environment interaction steps for which the initial base policy $\theta_{w_{i}}$ is trained under its weight $w_{i}$.
  • Retraining Preference Shift Strategy (controlled by shift magnitude $\delta_{s}$): Target scalarization weights $\{w_{i}^{\prime}\}$ for directional retraining are generated by shifting each initial weight $w_{i}$ to a nearby, distinct point on the preference space. The extent of this shift is controlled by a hyperparameter $\delta_{s}$. Conceptually, for $d$-dimensional preference spaces ($d>2$), this shift could be defined as an angular displacement in the space. In our current two-dimensional objective experiments ($d=2$), where $w_{i}=[w_{i,0},w_{i,1}]$, this shift is implemented by moving the first component $w_{i,0}$ by the magnitude $\delta_{s}$ to obtain $w^{\prime}_{i,0}$. The shift direction (decrease or increase) is chosen to keep the component within valid bounds (e.g., $[0,1]$), and the default direction is decrease; then the second component $w_{i,1}$is adjusted accordingly (assuming all objective weights sum to 1).
  • Directional Retraining Timesteps ($T_{\mathrm{dir}}$): The number of environment interaction steps for which the base policy $\theta_{w_{i}}$ is retrained under its target weight $w_{i}^{\prime}$ to produce $\theta_{w_{i}^{\prime}}$.
  • Step-Scale Factor Generation ($\alpha_{\text{start}},\alpha_{\text{end}},\Delta\alpha$): The set of step-scale factors $\{\alpha_{j}\}$ used in Locally Linear Extension is generated based on a starting value ($\alpha_{\text{start}}$), an ending value ($\alpha_{\text{end}}$), and either a step increment ($\Delta\alpha$).
  • Fine-tuning Timesteps ($T_{\mathrm{ref}}$): The number of environment interaction steps for which the selected candidate policy from the extension phase is fine-tuned under its matched preference weight $w_{i,j}$.
The specific values for these LLE-MORL parameters, are provided in Table 4 and Table 5.
Parameter Name MO-Swimmer MO-Hopper-2d MO-Ant-2d
steps per actor batch 512 512 512
learning rate (×$\times 10^{-4}$) 3 3 3
learning rate decay ratio 1 1 1
$\gamma$ 0.995 0.995 0.995
GAE lambda 0.95 0.95 0.95
number of mini batches 32 32 32
PPO epochs 10 10 10
entropy coefficient 0.0 0.0 0.0
value loss coefficient 0.5 0.5 0.5
maximum gradient norm 0.5 0.5 0.5
clip parameter 0.2 0.2 0.2
@@
Parameter Name Symbol MO-Swimmer MO-Hopper-2d MO-Ant-2d
Number of base policies $K$ 6 6 6
Initialization timesteps $T_{\mathrm{init}}$ $1\times 10^{5}$ $1\times 10^{5}$ $1\times 10^{5}$
Preference shift magnitude $\delta_{s}$ 0.1 0.1 0.1
Directional retraining timesteps $T_{\mathrm{dir}}$ $1\times 10^{4}$ $1\times 10^{4}$ $1\times 10^{4}$
Step-scale start $\alpha_{\mathrm{start}}$1.5 -1.5 -1.5
Step-scale end $\alpha_{\mathrm{end}}$ 1.5 1.5 1.5
Step-scale increment $\Delta\alpha$ 0.05 0.05 0.05
Fine-tuning timesteps $T_{\mathrm{ref}}$ $1\times 10^{3}$ $1\times 10^{3}$ $1\times 10^{3}$
Table 4: Hyperparameters for the LLE-MORL across benchmarks under sample-efficient setting.
Parameter Name Symbol MO-Swimmer MO-Hopper-2d MO-Ant-2d
Number of base policies $K$ 6 6 6
Initialization timesteps $T_{\mathrm{init}}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$
Preference shift magnitude $\delta_{s}$ 0.1 0.1 0.1
Directional retraining timesteps $T_{\mathrm{dir}}$ $1\times 10^{4}$ $1\times 10^{4}$ $1\times 10^{4}$
Step-scale start $\alpha_{\mathrm{start}}$1.5 -1.5 -1.5
Step-scale end $\alpha_{\mathrm{end}}$ 1.5 1.5 1.5
Step-scale increment $\Delta\alpha$ 0.05 0.05 0.05
Fine-tuning timesteps $T_{\mathrm{ref}}$ $1\times 10^{3}$ $1\times 10^{3}$ $1\times 10^{3}$
Table 5: Hyperparameters for the LLE-MORL across benchmarks under standard-training setting.
B.3 Experiments compute resources
@@
All experiments were run on a workstation equipped with an AMD Ryzen Threadripper PRO 5975WX (32 cores), an NVIDIA GeForce RTX 3090 GPU (24 GB GDDR6X), and 256 GiB of RAM, running Ubuntu 24.04 LTS. The software stack included CUDA Toolkit 12.0 and the corresponding NVIDIA drivers. Approximate execution times for all methods and benchmarks are reported separately in Table 6.
Sample-efficient setting Standard-training setting
Method MO-Swimmer MO-Hopper-2d MO-Ant-2d MO-Swimmer MO-Hopper-2d MO-Ant-2d
GPI-LS 3 3 3 22 23 19
CAPQL 13 8 12 95 77 68
MORL/D 1 1 1 3 3 3
LLE-MORL 1 1 1 3 3 3
@@
Table 6: Approximate execution times (hours) for each method and benchmark under sample-efficient and standard training setting.
Appendix C Experiment Statistical Significance
Our demonstration of reliability and robustness relies on: (1) the consistent observation of the superior performance of LLE-MORL in generating high-quality Pareto fronts compared to baselines; (2) sensitivity analyses of key hyperparameters (detailed in Appendix E), which show consistent outcome patterns within certain parameter ranges; and (3) the qualitative consistency in the shape, extent, and dominance characteristics of the visualized Pareto fronts (e.g. Figure 4 and Figure 5). Although this study prioritises the introduction of our novel mechanism over exhaustive multi-seed statistical tests for all reported metrics, these combined observations provide confidence in the principal findings and effectiveness of LLE-MORL.
Appendix D Limitations
  • We have restricted ourselves to problems with two objectives where the Pareto front is one-dimensional. A larger number of objectives is a problem for most of the existing MORL algorithms. Although usually some of the objectives are of different importance and can be lexicographically ranked, so that the complexity does not necessarily increase exponentially with the number of objectives. The high-dimensional case is nevertheless challenging, but our approach can be seen as promising: In higher dimensions, the number of solutions that are in the same local quasi-linear patch increases dramatically, so that the efficiency of the proposed local search will be even more beneficial. This benefit could be reduced by the potentially increasing complexity of the topological relation between the performance space and the parameter space which could be a fascinating subject for future work.
  • We are assuming that the Pareto front consists of a relatively small number of connectivity components which have a manifold structure. While there is no theoretical bound to the complexity of the Pareto from, the idea of MORL implies that the objectives are at least in some sense comparable. For Pareto fronts that are fractal or of high genus, the result of multi-objective optimisation lacks robustness, although it will neither be possible to fix any limits for the complexity of the Pareto front. However, as long as there are only a limited number of manifold-like connectivity components, our algorithm will be applicable.
  • Widely different scales and elasticities of the objectives can lead to problems as in optimisation in anisotropic error landscapes. Step size control that helps in gradient methods in optimisation, will also be useful here, but has not been studied yet, as the typical (benchmark) problems are sufficiently isotropic.
  • The density of the identified solutions on the Pareto front is clearly a challenge which may be solved by step size control as mentioned in the previous point. This concerns higher-dimensional cases as well as extended one-dimensional trails as visible in the top trail in Figure 3c also a simple reduction of the parameter $\Delta\alpha$ at the observation of large steps in the performance space could have solved this issue already so that a more uniform covering of the Pareto from is not difficult to achieve in the present approach. See also App. E.2. In contrast to other approaches, linear regions of the Pareto can trivially be tracked by LLE-MORL. Concave regions connected to the Pareto front will be followed through without problem, but will need to be removed in a single postprocessing step as they are dominated by other solutions. Even full patches of solutions may turn out to be Pareto sub-optimal and require a similar treatment.
  • We have made use of scalarisation to seed the solution domains, whereas the reconstruction of the Pareto front is done by a lateral process that does not depend on preference weights. It is in principle possible that a solution patch is not reachable by any scalarisation-based seeding attempt, see also the early discussion in [Vamplew et al., 2008]. In this case our approach might not find this patch, although it is still possible that it is found by retraining from a different solution domain as shown in Figure 3c.
Appendix E Additional Results
E.1 Effect of Directional Retraining Shift

### v2

@@ C.1 Benchmarks
To evaluate the performance of our proposed LLE-MORL method and compare it against existing baselines, we utilise a suite of continuous control benchmarks from the MO-Gymnasium library (Felten et al., 2023). These environments are designed to test the ability of an agent to learn policies that effectively balance multiple, often conflicting objectives. The specific environments and their multi-objective reward formulations are detailed below:
@@
A planar, three-link swimmer operating in a viscous fluid, utilising a 2D continuous action space to control its joint torques. The objectives are to maximise forward velocity along the $x$-axis and minimise the control cost.
The observation space $\mathcal{S}\subset\mathbb{R}^{8}$ includes joint angles and velocities, and the action space $\mathcal{A}\subset\mathbb{R}^{2}$ represents joint torques in [1,1][-1,1]. Let $x_{\text{before}}$ and $x_{\text{after}}$ be the $x$-coordinates of the centre of mass of swimmer before and after an action, $\Delta t$ be the time step, and $a_{j}$ be the $j$-th component of the action vector.
@@
This environment features a 2D one-legged hopper with a 3-dimensional continuous action space controlling torques for its thigh, leg, and foot joints. Originally a 3-objective task (forward speed, jump height, control cost), we use the 2-objective variant, in which the separate control-cost objective is added to other objectives.
The observation space $\mathcal{S}\subset\mathbb{R}^{11}$ includes joint states and torso position, and the action space $\mathcal{A}\subset\mathbb{R}^{3}$ represents joint torques in [1,1][-1,1]. Let $v_{x}=(x_{\text{after}}-x_{\text{before}})/{\Delta t}$ be the forward velocity of the agent along the $x$-axis, where $x_{\text{after}}$ and $x_{\text{before}}$ are x-positions of the torso. Let $h_{\text{jump}}=10\times(z_{\text{after}}-z_{\text{init}})$ be a measure of jumping height, where $z_{\text{after}}$ is the current z-position of the torso and $z_{\text{init}}$ is its initial z-position. Let $c_{\text{ctrl}}$ be the positive control cost, computed as $w_{\text{env\_ctrl}}\sum_{j}(a_{j})^{2}$, where $w_{\text{env\_ctrl}}$ is the environment control cost weight (typically $0.001$). Let $r_{\text{healthy}}$ be the health reward (typically +1+1 if the agent has not fallen). The reward vector $\mathbf{R}=[R_{1},R_{2}]$ is defined as:
@@
A quadrupedal “ant” robot in 2D with an eight-dimensional action space for joint torques. By default, the environment emits a three-dimensional reward vector: (1) $x$-velocity, (2) $y$-velocity, and (3) control cost. Here, we use the two-objective variant in which the separate control-cost objective is added to other objectives.
The observation space $\mathcal{S}\subset\mathbb{R}^{27}$ includes joint states, torso position, and contact forces, and the action space $\mathcal{A}\subset\mathbb{R}^{8}$ represents joint torques in [1,1][-1,1]. Let $v_{x}=(x_{\text{after}}-x_{\text{before}})/{\Delta t}$ be the forward velocity of the agent along the $x$-axis, where $x_{\text{after}}$ and $x_{\text{before}}$ are x-positions of the torso. Let $v_{y}=(y_{\text{after}}-y_{\text{before}})/{\Delta t}$ be the forward velocity of the agent along the $y$-axis, where $y_{\text{after}}$ and $y_{\text{before}}$ are y-positions of the torso. Let $c_{\text{ctrl}}$ be the positive control cost, computed as $w_{\text{env\_ctrl}}\sum_{j}(a_{j})^{2}$, where $w_{\text{env\_ctrl}}$ is the environment control cost weight (typically 0.05). Let $r_{\text{healthy}}$ be the health reward (typically +1+1 if the Ant is healthy). Let $p_{\text{contact}}$ be the positive contact penalty, which is used for penalising the Ant if the external contact forces are too large, computed as $w_{\text{env\_contact}}\sum_{k}(\text{force}_{k})^{2}$, where $w_{\text{env\_contact}}$ is the environment contact cost weight (typically $5\times 10^{-4}$). The reward vector $\mathbf{R}=[R_{1},R_{2}]$ is defined as:
@@
MO-Hopper-3obj-v5.
This environment is identical to MO-Hopper-2obj-v5 in terms of state and action spaces, but retains all three original objectives instead of merging the control cost into the other rewards. As a result, the task features three objectives: forward speed, jump height, and control cost. The reward vector $\mathbf{R}=[R_{1},R_{2},R_{3}]$ is defined as:
  • $R_{1}$ ($x$-Velocity Performance):
    $R_{1}=v_{x}+r_{\text{healthy}}$
  • $R_{2}$ ($y$-Velocity Performance):
    $R_{2}=h_{\text{jump}}+r_{\text{healthy}}$
  • $R_{3}$ (Control Cost):
    $R_{3}=-\,c_{\text{ctrl}}+r_{\text{healthy}}$
MO-Ant-3obj-v5.
This environment is identical to MO-Ant-2obj-v5 in terms of state and action spaces, but retains all three original objectives instead of merging the control cost into the other rewards. As a result, the task features three objectives: $x$-velocity, $y$-velocity, and control cost. The reward vector $\mathbf{R}=[R_{1},R_{2},R_{3}]$ is defined as:
  • $R_{1}$ ($x$-Velocity Performance):
    $R_{1}=v_{x}+r_{\text{healthy}}-p_{\text{contact}}$
  • $R_{2}$ ($y$-Velocity Performance):
    $R_{2}=v_{y}+r_{\text{healthy}}-p_{\text{contact}}$
  • $R_{3}$ (Control Cost):
    $R_{3}=-\,c_{\text{ctrl}}+r_{\text{healthy}}-p_{\text{contact}}$
C.2 Evaluation Metrics
Hypervolume (HV). Let $P$ be an approximate Pareto front and $G_{0}$ a reference point dominated by all $p\in P$. The hypervolume is $\mathcal{H}(P)=\int_{\mathbb{R}^{d}}\mathbbm{1}_{H(P)}(z)\,dz,$ where $H(P)=\{\,z\in\mathbb{R}^{n}\mid\exists\,j,\;1\leq j\leq|P|\;:\;G_{0}\preceq z\preceq P(j)\}$. Here, $P(j)$ is the $j^{\rm th}$ solution in $P$, the symbol $\preceq$ denotes objective dominance, and $\mathbbm{1}_{H(P)}$ is an indicator function that equals 1 if $z\in H(P)$ and 0 otherwise. A higher hypervolume implies a front closer to and more extensive with respect to the true Pareto front.
C.3 Hungarian matching distance
To measure structural similarity between policies, we use the Hungarian matching distance (Kuhn, 1955; Munkres, 1957).
Definition C.1.
Hungarian matching distance. For a given layer $l$ with neuron sets $A^{(l)}$ and $B^{(l)}$, let $w_{i}^{(l)}$ and $w_{j}^{(l)}$ denote the incoming weight vectors of neurons $i\in A^{(l)}$ and $j\in B^{(l)}$, respectively. The minimum-cost perfect matching $M^{(l)}$ between $A^{(l)}$ and $B^{(l)}$ is obtained by using the Hungarian algorithm:
$M^{(l)}=\underset{\text{matching}}{\arg\min}\sum_{(i,j)\in\text{matching}}\left\lVert w_{i}^{(l)}-w_{j}^{(l)}\right\rVert_{2}$ (43)
The Hungarian matching distance between the two networks is then defined as
$d_{\text{Hungarian}}(A,B)=\sum_{l=1}^{L}\sum_{(i,j)\in M^{(l)}}\left\lVert w_{i}^{(l)}-w_{j}^{(l)}\right\rVert_{2}$ (44)
where $L$ is the total number of layers.
C.4 Sanity Check Details
In Section 3.2, the sanity check was designed as a qualitative validation of the parameter–performance relationship (PPR), rather than a full experiment.
To construct policy pairs, we trained 11 base policies $\theta_{\omega 1}$ using the scalarization vector $\omega_{1}$ that were equally distributed in the preference space, i.e. $\omega_{1}=[1,0],[0.9,0.1],...,[0,1]$.
For each base policy, we formed a paired scalarization vector $\omega_{2}$ by shifting $\omega_{1}$ with a retraining shift parameter $\delta_{s}$:
$\omega_{2}=\begin{cases}[\omega_{11}-\delta_{s},\omega_{12}+\delta_{s}],\text{ if }\omega_{11}-\delta_{s}\in[0,1]\\ \\ [\omega_{11}+\delta_{s},\omega_{12}-\delta_{s}],\text{ otherwise}\end{cases}$ (45)
where $\delta_{s}\in\{0.1,0.2,0.3,0.4,0.5\}$. This procedure yielded multiple pairs of $(\omega_{1},\omega_{2})$ across the preference space. For each pair, we performed short retraining from $\theta_{\omega_{1}}$ under $\omega_{2}$ and compared the orignial $\theta_{\omega_{1}}$ and resulting model $\theta_{\omega^{\prime}}$ with the independently trained $\theta_{\omega_{2}}$. This allowed us to consistently observe that retrained policies stay close in parameter space (low Hungarian distance), while having directional movement in performance space, supporting the local PPR hypothesis. Figure 2 in the main paper presents a representative example from these trials.
C.5 Training Details
All learning phases within our LLE-MORL algorithm, including the initial training of base policies, the directional retraining, and the final preference-aligned fine-tuning, utilize the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017). We employed a standard PPO implementation from the Stable Baselines3 library (Raffin et al., 2019). The PPO parameters used across all training stages and benchmarks are detailed in Table 3.
@@ C.5 Training Details
  • Number of Initial Base Policies ($K$): The total count of base policies $\theta_{w_{k}}$, trained in the initialization stage. The corresponding $K$ initial scalarization weights $\{w_{k}\}_{k=1}^{K}$ are generated by evenly distributing them across the preference space (e.g., for 2D objectives, from $[1,0]$ to $[0,1]$ in $K$ steps).
  • Initialization Training Timesteps ($T_{\mathrm{init}}$): The number of environment interaction steps for which the initial base policy $\theta_{w_{k}}$ is trained under its weight $w_{k}$.
  • Retraining Preference Shift Strategy (controlled by shift magnitude $\delta_{s}$): Target scalarization weights $\{w_{k}^{(i)}\}$ for directional retraining are generated by shifting each initial weight $w_{k}$ to a nearby, distinct point on the preference space. The extent of this shift is controlled by a hyperparameter $\delta_{s}$.
  • Directional Retraining Timesteps ($T_{\mathrm{dir}}$): The number of environment interaction steps for which the base policy $\theta_{w_{k}}$ is retrained under its target weight $w_{k}^{i}$ to produce $\theta_{w_{k}^{(i)}}$.
  • Step-Scale Factor Generation ($\alpha_{\text{start}},\alpha_{\text{end}},\Delta\alpha$): The set of step-scale factors $\{\alpha_{i}\}$ used in Locally Linear Extension is generated based on a starting value ($\alpha_{\text{start}}$), an ending value ($\alpha_{\text{end}}$), and either a step increment ($\Delta\alpha$).
  • Fine-tuning Timesteps ($T_{\mathrm{ref}}$): The number of environment interaction steps for which the selected candidate policy from the extension phase is fine-tuned under its matched preference weight $w_{k,\text{cand}}$.
@@
Parameter Name MO-Swimmer MO-Hopper-2d MO-Ant-2d MO-Hopper-3d MO-Ant-3d
steps per actor batch 512 512 512 512 512
learning rate ($\times 10^{-4}$) 3 3 3 3 3
learning rate decay ratio 1 1 1 1 1
$\gamma$ 0.995 0.995 0.995 0.995 0.995
GAE lambda 0.95 0.95 0.95 0.95 0.95
number of mini batches 32 32 32 32 32
PPO epochs 10 10 10 10 10
entropy coefficient 0.0 0.0 0.0 0.0 0.0
value loss coefficient 0.5 0.5 0.5 0.5 0.5
maximum gradient norm 0.5 0.5 0.5 0.5 0.5
clip parameter 0.2 0.2 0.2 0.2 0.2
In all experiments reported in the main results, we fix the number of initial base policies to $K=6$, the retraining preference shift magnitude to $\delta_{s}=0.1$, and the step-scale factor grid to $\alpha\in[-1.5,1.5]$ with step size $\Delta\alpha=0.05$. We need to clarify that $K$ and $\delta_{s}$ are chosen for simplicity and are not intended to be optimal. We have an additional discussion about the choice of $\delta_{s}$ in Appendix D.1. The step-scale factor $\alpha$ controls the Locally Linear Extension stage and has a meaningful but limited influence on performance; its impact and selection are discussed in Appendix D.2.
For baseline methods, since the original papers do not evaluate performance on the same benchmark problems considered in this work, we adopt the implementations and hyperparameter settings provided by morl-baselines (Felten et al., 2023). As LLE-MORL is a PPO-based algorithm, we employ PPO hyperparameter configurations commonly used in PPO-based MORL methods, such as PGMORL (Xu et al., 2020), which serves as a standard and widely adopted reference for PPO settings in MORL tasks.
In all experiments reported in the main results, we strictly control the total training interaction budget for all methods to ensure a fair comparison. For LLE-MORL, this fixed budget is automatically partitioned across the three learning phases using a simple 3:1:13\!:\!1\!:\!1 ratio for initial base policy training, directional retraining, and preference-aligned fine-tuning, respectively. When the total training budget is sufficiently large, allocating additional interaction steps to the base policy training stage can further improve performance, as directional retraining and fine-tuning require fewer updates. The optimal budget allocation depends on the target application and can be adjusted on a case-by-case basis.
C.6 Computational Resources
@@
All experiments were run on a workstation equipped with an AMD Ryzen Threadripper PRO 5975WX (32 cores), an NVIDIA GeForce RTX 3090 GPU (24 GB GDDR6X), and 256 GiB of RAM, running Ubuntu 24.04 LTS. The software stack included CUDA Toolkit 12.0 and the corresponding NVIDIA drivers. Approximate execution times for all methods and benchmarks are reported separately in Table 4(c).
@@
Table 4: Approximate execution times (hours) for each method and benchmark under different settings.
(a) 2D sample-efficient setting
Method MO-Swimmer MO-Hopper-2d MO-Ant-2d
GPI-LS 3 3 3
CAPQL 3 3 3
Q-Pensieve 3 3 3
MORL/D 1 1 1
LLE-MORL 1 1 1
(b) 2D standard-training setting
Method MO-Swimmer MO-Hopper-2d MO-Ant-2d
GPI-LS 22 23 19
CAPQL 11 11 12
Q-Pensieve 17 15 18
MORL/D 3 3 3
LLE-MORL 3 3 3
(c) 3D setting
Method MO-Hopper-3d MO-Ant-3d
GPI-LS 44 43
CAPQL 13 18
Q-Pensieve 24 23
MORL/D 5 6
LLE-MORL 5 5
Appendix D Additional Results
D.1 Effect of Directional Retraining Shift